OpenStudy (anonymous):

16/28 in lowest terms

OpenStudy (anonymous):

\[\frac{ 16 }{ 28 }\] Divide both numbers by a number they both can go into evenly, in this case 2 or 4

OpenStudy (calculusxy):

1) find the GCF (greatest common factor) - 4 2) divide 16 and 28 by 4 - 16/4=4 and 28/4=7 Answer-4/7

OpenStudy (anonymous):

16/28=8/14= 4/7

OpenStudy (anonymous):

8/9

OpenStudy (anonymous):

thanks everyone

OpenStudy (calculusxy):

No prob

OpenStudy (pixiedust1):

Do you understand the concept now?